OpenStudy (anonymous):

How do you factor these perfect square trinomials step by step?: x^2-16x+256 4x^2-4x+1

OpenStudy (anonymous):

the first one is not a perfect square?

OpenStudy (anonymous):

Would it be considered a prime?

OpenStudy (anonymous):

it is prime since it can be factored

OpenStudy (anonymous):

What about the second one?

OpenStudy (anonymous):

it is a perfect square trinomial,(2x-1)^2..

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

you surely know the formulas:(a+b)^2=a^2+2ab+b^2 and (a-b)^2=a^2-2ab+b^2 in your example,you must guess which term when squared gives 4x^2 which is 2x which is now a and which number when squared gives 1 which is 1,your b now you still have to check if -2ab=-4x -2(2x)(1)=-4x so its (2x-1)^2